OpenStudy (anonymous):

help me, I dont understand!!!! 4j^2k^6(2j^11)^3k^5

OpenStudy (anonymous):

its ambiguous please use more brackets or the equation button

OpenStudy (anonymous):

ummmm, ok

OpenStudy (anonymous):

\[4j ^{2}k ^{6}(2j ^{11})^{3}k ^{5}\]

OpenStudy (anonymous):

\[x^a * x^b = x^{a+b}\]\[(x^a)^b = x^{ab} = (x^b)^a\]

OpenStudy (anonymous):

Im not sure I understand!! :\

OpenStudy (anonymous):

so we have\[4j^2* k^6* k^5*(2j^{11})^3 = 4j^2 * k^{6 + 5}* 2^3 * j^{11*3} = 4j^2 * k^{11}* 8*j^{33}\]\[=32j^{35}k^{11}\]

OpenStudy (anonymous):

Thank you @Euler271

OpenStudy (anonymous):

happy to help :)